an inlet pump can fill a water tower in 15 hours and an outlet pump can empty the tower in 20 hours. if both pumps are left open, how long will it take the tower to fill up?
Answer by CharlesG2(834)

an inlet pump can fill a water tower in 15 hours and an outlet pump can empty the tower in 20 hours. if both pumps are left open, how long will it take the tower to fill up?
inlet fills 1/15 tower in an hour
outlet empties 1/20 tower in an hour
every hour tower fills 1/15 - 1/20 = 4/60 - 3/60 = 1/60
so 60 hours to fill tower